ABC, Inc. is looking at raising additional capital for the future project. The project is expected to provide a return on investment of 13%. In order for ABC, Inc. to determine whether this project is worth investing in, it must first determine the cost of capital it will use to finance the project. a. The firm's current stock price is $45 and it has 4 million shares of stock outstanding. The firm also has $30 million of preferred stock and $70 million of debt. Calculate the weights of each capital component. b. The firm is looking at issuing a new 30-year bond that pays an annual coupon of 8% with a flotation cost of 2%. The bond is expected to sell at its par value of $1000. The firm's tax rate is 40%. Calculate the ATrd. c. The firm will have to pay the underwriter a 10% flotation cost for the new equity it will raise. The firm just paid out a dividend of $4.22 with an expected growth rate of 4.5%. Calculate the re d. The firm expects its preferred stocks to sell for $112.55. The par value of the preferred stock will be $100 with a 12% annual dividend. The flotation cost will be paid to the underwriter will be 4%. Calculate the rps. e. Assume that the firm's current market value is their target capital structure, what is the firm's WACC? f. Should the firm take on this investment based on the cost of the capital that it will use to fund the project? Why?
Answer #1

All financials below are in $ mn.

a. The firm's current stock price is $45 and it has 4 million shares of stock outstanding. The firm also has $30 million of preferred stock and $70 million of debt. Calculate the weights of each capital component.

Common equity, E = Price x nos. of shares outstanding = 45 x 4 = 180

Preferred stock, P = 30

D = 70

Total capital, T = E + P + D = 180 + 30 + 70 = 280

Weight of equity, We = E / T = 180 / 280 = 64.29%

Weight of preferred stock, Wps = P / T = 30 / 280 = 10.71%

Weight of debt, Wd = D / T = 70 / 280 = 25.00%

b. The firm is looking at issuing a new 30-year bond that pays an annual coupon of 8% with a flotation cost of 2%. The bond is expected to sell at its par value of $1000. The firm's tax rate is 40%. Calculate the ATrd.

We can find the YTM of the bond using RATE function in excel. Inputs are:

Period = 30, PMT = payment per period = annual coupons = 8% x Par value = 8% x 1,000 = 80, PV = - Price of bond = - Par value x (1 - Flotation cost) = - 1,000 x (1 - 2%) = -980; FV = future value = par value = -1000

Hence, YTM = RATE ( Period, PMT, PV, FV) = RATE (30, 80, -980, 1000) = 8.18%

After tax cost of debt, ATrd = YTM x (1 - T) = 8.18% x (1 - 40%) = 4.91%

c. The firm will have to pay the underwriter a 10% flotation cost for the new equity it will raise. The firm just paid out a dividend of $4.22 with an expected growth rate of 4.5%. Calculate the re

D0 = 4.22; g = 4.5%; Current share price, S = 45; Flotation expenses, F = 10%

D1 = D0 x (1 + g) = 4.22 x (1 + 4.5%) = 4.41

Cost of equity, re = D1 / [S x (1 - F)] + g = 4.41 / [45 x (1 - 10%)] + 4.5% = 15.39%

d. The firm expects its preferred stocks to sell for $112.55. The par value of the preferred stock will be $100 with a 12% annual dividend. The flotation cost will be paid to the underwriter will be 4%. Calculate the rps.

Annual dividend, D = 12% x Par value = 12% x 100 = 12

Price, P = 112.55; Flotation, F = 4%

Hence, cost of preferred stock = rps = D / [P x (1 - F)] = 12 / [112.55 x (1 - 4%)] = 11.11%

e. Assume that the firm's current market value is their target capital structure, what is the firm's WACC?

WACC = Wd x ATrd + Wps x rps + We x re = 25% x 4.91% + 10.71% x 11.11% + 64.29% x 15.39% = 12.31%

f. Should the firm take on this investment based on the cost of the capital that it will use to fund the project? Why?

ROI = 13% > WACC = 12.31%

Hence, the firm should take on this investment.